Solve the following recurrence relations and give the value of f(N)

f(n) = -1 for n= 0

f(n) = f(n-1)+ n for n>0
Answer #1

f(n)=-) f(n=f(n-1) + o 2 = f(n= fcirati i fronti salt f(2)= f (2-1) +2 f(12+2 o+2l=2 f (3-1)+3 = f(z)t3 =2+3 f (4)= f (4-10 +